Question 2                                                                                                           (Total: 40 marks)

Tabulation Corporation manufactures and sells two types of electronic calculators: EL-520 W and EL-620 T. The following data was gathered from last month’s activities:

                                                                                                             EL-520 W             EL-620 T

Sales in units                                                                                   5,000                     3,000

Selling price per unit                                                                    $50                         $100

Variable production costs per unit                                         $10                         $26

Traceable fixed production costs                                            $100,000              $150,000

Variable selling expenses per unit                                          $5                           $6

Traceable fixed selling expenses                                            $5,000                   $7,500

Allocated division administrative expenses                        $50,000                 $60,000

Required:

  1. Prepare a segmented income statement in the contribution format for last month, showing both "Amount" and "Percent" columns for the company as a whole and for each model.
  2. Why might it be very difficult to calculate separate break-even sales for each model?
  3. Refer to the original data and, if necessary, the results of the segmented income statement prepared in part (1) above. Calculate the total break-even sales (in both units AND dollars) for last month, assuming that none of the fixed production costs and fixed selling expenses is traceable. Allocate the total break-even sales between the two models.
  4. Again, refer to the original data and, if necessary, the results of the segmented income statement prepared in part (1) above. Calculate the total break-even sales (in both units AND dollars) for last month, assuming that the "allocated" amounts of the company's administrative expenses are actually traceable. Allocate the total break-even sales between the two models.
  5. How reasonable are the total break-even sales numbers calculated in parts (3) and (4) given the actual results for last month?
Answer #1

ANSWER

AS per the given question.

Note:Only 4 sub parts are allowed to be answered, kindly comment in case of any doubt or difficulty.

Answer 1.

Segments Total Company EL-520 W EL-620 T $550,000 $250,000 $300,000 100% Sales 100% 100% Variable expenses $171,000 $75000 $9

Answer 2.

It is difficult to calculate break even analysis because fixed costs per unit for each model is not traceable and allocated.

Answer 3.

Break even in units = Fixed costs / contribution margin per unit

Break even in units = (262500 + 110000) / (379000 / 8000) =7862.7 = 7863 units

Break even in sales = fixed costs / cm ratio = (262500 + 110000) / 0.69  = 539855 (it can also be 540560 or 540567, depending upon how much we round off)

No allocation is possible since we are assuming fixed costs are not traceable.

Answer 4.

we are assuming fixed costs are traceable.

EL-520 W  

Break even in units = 155000 / 35 = 4428.57= 4429 units

Break even in sales = 155000 / 0.70 = 221,429 (or 4429*50 = 221450 if we follow this formula)

EL-620 T

Break even in units = 217500 / 68 = 3198.529 = 3199 units

Break even in sales = 217500 / 0.68 = 319853 (or 3199*100 = 319900 if we follow this formula)

Total Break even in units = 3199 + 4429 = 7628 units

Break even in sales = 319853 + 221,429 = 541282

KINDLY RATE